### v2

# The complete action for $\mathcal{N}=2$ de Sitter pure supergravity

Abstract
Supergravity theories in de Sitter spacetime are known to be very constrained, and rather unnatural within String/M Theory. We revisit the seminal paper by Pilch, van Nieuwenhuizen and Sohnius, where the possible existence of a real Lagrangian for ${\cal N}=2$ pure supergravity in four-dimensional de Sitter spacetime was pointed out. We clarify several issues related to the non-unitarity of the theory and explicitly construct the unique, complete theory searched for long ago by the aforementioned authors. We argue that the lack of unitarity of the Lorentzian theory may be revisited in the Euclidean approach to de Sitter quantum gravity, where alternative definitions of unitarity can be introduced.
1 Introduction
On the one hand, both the primordial and current Universe seem to be well approximated by de Sitter geometry. On the other hand, the non-renormalisability of Einstein’s theory of gravity calls for an extension of it that should behave better in the high-energy limit. Supergravities were born in the late seventies [1, 2] with the hope to give a renormalisable model of quantum field theory containing Einstein’s theory as a consistent truncation. This hope was later shown to be unrealistic, string theory offering more promises to attain UV finiteness while at the same time containing a massless spin-2 mode associated with the graviton, see e.g. [3]. More recently, however, the works [4, 5, 6, 7, 8, 9] and references therein prompted a renewed interest in supergravity theories. The very first paper where supergravity was considered in a de Sitter background is [10] by Pilch, van Nieuwenhuizen, and Sohnius. At that time, when supergravity theories were born, our Universe was believed to have no cosmological constant, and the paper [10] appeared to be of more academic relevance than phenomenological necessity. Still, this paper is seminal in that, for the first time, it clearly described the problems encountered in constructing supergravity theories in de Sitter (dS4) spacetime due to a clash between the reality conditions and the necessary gauge invariance to be imposed on a vector-spinor to only propagate the helicity degrees of freedom $\pm 3/2$. Supergravity favours a negative cosmological constant [11] — although this was not stressed in the seminal work [11], which also adopted misleading terminology in use at the time, when the maximally symmetric spacetime of negative constant curvature was termed “de Sitter” instead of “anti-de Sitter” —, therefore it was crucial that the paper [10] pointed out the possible existence of a consistent supergravity theory in de Sitter spacetime at the price of a ghostly graviphoton, also showing that the minimal number of supersymmetries in dS4 supergravity is ${\cal N}=2\,$.
Since then, it was discovered [12, 13, 14] that our Universe possesses a small, positive cosmological constant, de Sitter spacetime being the simplest and most symmetric solution to Einstein’s equations with a positive cosmological constant. String/M theory, although remarkable in that it is a quantum theory that contains gravity, does not seem to naturally allow for de Sitter vacua, see e.g. [15, 16, 17], but see also [18, 19, 20] for some positive results in this direction. Independently of the issue of de Sitter vacuum in string theory, the study of quantum field theories in de Sitter spacetime is a very important topic, dating back to the sixties, see for example [21, 22, 23, 24, 25, 26] for a very small set of references. The field has witnessed a recent surge of interest, see e.g. [27, 28, 29, 30, 31] and references therein.
In the present paper, we revisit the work [10] and complete it in several directions. First, we rigorously prove the existence and uniqueness of the simplest possible supergravity theory in four-dimensional de Sitter (dS4) spacetime, using tools [32, 33] unavailable at the time of [10]. The latter work gives an action invariant under gauge supersymmetry, only up to terms quartic in fermions, while in the present paper we provide the unique, complete Lagrangian that is invariant under ${\cal N}=2$ local supersymmetry, to all orders in the coupling constant (Newton’s constant), and in particular, to all orders in the fermionic fields. Second, we provide insight concerning the issue of non-unitarity discussed in [10]. In particular, although we confirm that the graviphoton necessarily has the wrong sign in front of its kinetic term, we explain that the massless gravitino also has unitarity issues in de Sitter, not observed at the time of [10]. Using recent results [29, 34, 35] showing that the quantisation of the free Dirac gravitino on Lorentzian dS4 leads to indefinite-norm states, we clarify that the graviphoton is not the only source of non-unitarity for $\mathcal{N}=2$ de Sitter supergravity, contrary to the standard lore.
We conclude this paper by discussing the relevance of our results in the context of the Euclidean approach to quantum gravity with positive cosmological constant, suggesting future directions. The Euclidean approach to de Sitter quantum gravity enjoys renewed interest thanks to invigorating recent results [36, 37, 38], such as 1-loop expressions for the sphere path integral in terms of group-theoretic characters of the de Sitter group. We explain how our results for the existence and uniqueness of pure $\mathcal{N}=2$ de Sitter supergravity in Lorentzian signature set the basis for the study of the Euclidean version of the theory. Motivated by recent findings for the 1-loop sphere partition function for the gravitino [38], we explain why the unitarity issues of the Lorentzian supergravity theory may turn out to be harmless in Euclidean signature, suggesting promising future directions. In particular, this invites us to emphasise the importance of addressing a particular open question; the question of consistency of pure $\mathcal{N}=2$ de Sitter supergravity in Euclidean signature, see also discussions in [38].
Let us now review what is known for the two kinds of obstacles to the existence of unitary, unbroken de Sitter supersymmetry. [^1] For references about broken $\mathcal{N}=1$ supersymmetry and de Sitter, see e.g. [39, 40, 41, 42, 43, 44].
  1. 1.
    Non-unitarity of global supersymmetry on a fixed de Sitter background. This problem stems from abstract representation theory of the de Sitter superalgebra [10, 45]. The de Sitter superalgebra is the supersymmetric extension of the dS algebra $so(4,1)$, i.e. apart from the ten Grassmann-even $so(4,1)$ generators, there also exist spinorial supercharges ${Q}^{(i)}_{A}$ (Grassmann-odd generators). The index $A$ is a spinor index referring to the fundamental spinor representation of $so(4,1)$, and $i$ is an extended supersymmetry index. Let us assume that the spinorial generators are Dirac spinors for simplicity - however, the analysis presented in this paragraph holds also for the case where a symplectic Majorana reality condition is introduced for the supercharges [10, 45]. The main feature of the dS superalgebra is that the following trace vanishes
    $\sum_{A,i}\{{Q}^{(i)}_{A},{Q}^{(i)A\,\dagger}\}=0\;.$ (1)
    This is physically justified by the absence of a generator of $so(4,1)$ with bounded spectrum that could be placed on the right-hand side of (1). The vanishing of the trace in (1) implies that, if one tries to realise these spinorial supercharges as operators acting on a Hilbert space, then all non-trivial representations of the dS superalgebra must have indefinite norm. Similar unitarity no-go theorems exist for the super-extensions of $so(3,1)$ and $so(5,1)$ [45]. Interestingly, unitary representations exist for the super-extension of $so(2,1)$ [45]. See [46] for a recent discussion of the Euclidean approach to 2-dimensional supergravity with positive cosmological constant, where unitarity obstacles can be avoided.
  2. 2.
    Non-unitarity of dS4 pure supergravity and lack of a full theory. As mentioned earlier, a local and real action functional for $\mathcal{N}=2$ supergravity with positive cosmological constant, which is invariant under local supersymmetry, has been proposed in [10], where the terms quartic in fermion were not given. The dynamical fields of this theory are: a real vierbein, a real (gravi)photon, and two symplectic Majorana gravitini. The double number of symplectic Majorana gravitini is related to the fact that a conventional Majorana condition is not available, see e.g. [10, 35]. More specifically, this difficulty stems from the fact that the supercharges of the de Sitter superalgebra must behave as spinors of $Spin(4,1)$, and there is no Majorana spinor representation for this group. Nevertheless, a symplectic Majorana condition can still be used. We stress that the terms that are quartic in gravitini were not provided in [10], i.e. the local supersymmetry invariance of the action was checked up to terms quartic in fermions. The main result of Ref. [10] was that, although the $\mathcal{N}=2$ supergravity action is invariant under local supersymmetry (up to terms quartic in fermions), the photon kinetic term has the wrong sign, leading to the presence of negative norms. Thus, if a full theory exists, then it will have negative-norm states. However, the questions of whether a full theory of $\mathcal{N}=2$ supergravity with positive cosmological constant exists, and whether the theory is unique, were left open in [10]. The present paper addresses these questions.
2 The free theory
Consider a free model on a dS4 background with metric $\bar{g}_{\mu\nu}$ and cosmological constant $\Lambda=3\lambda^{2}$ containing: a Dirac massless spin-3/2 vector-spinor field $\phi_{\mu}$ [spinor indices are suppressed], a real massless spin-2 symmetric tensor field $h_{\mu\nu}$, and a real massless spin-1 field $A_{\mu}$. The action reads
$S_{0}=\int d^{4}x\sqrt{-\bar{g}}($ $-\bar{\phi}_{\mu}\gamma^{5}\gamma^{\mu\nu\rho}\Phi_{\nu\rho}-\sigma\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$ (2)
$-\frac{1}{2}\nabla^{\rho}h^{\mu\nu}\nabla_{\rho}h_{\mu\nu}+\nabla_{\rho}h^{\mu\nu}\nabla_{\mu}h_{\nu}{}^{\rho}-\nabla_{\mu}h\nabla_{\nu}h^{\mu\nu}$
$+\frac{1}{2}\nabla^{\mu}h\nabla_{\mu}h+3\lambda^{2}h_{\mu\nu}h^{\mu\nu}-\frac{3}{2}\lambda^{2}h^{2})\;,$
where $h=h^{\mu}\,_{\mu}$, $\Phi_{\nu\rho}=\nabla_{[\nu}\phi_{\rho]}+\frac{i}{2}\lambda\gamma_{[\nu}\phi_{\rho]}$ and $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$  [^2] The symbol $\nabla_{\mu}$ denotes the Lorentz-covariant derivative of the background de Sitter geometry, where we use the convention $\nabla_{\mu},\nabla_{\nu}]\psi=\frac{1}{2}\lambda^{2}\gamma_{\mu\nu}\psi$, with $\psi$ a four component spinor. We follow the strength-one (anti)symmetrisation convention, with (square) round brackets indicating (anti)symmetrisation of the enclosed indices. For example, $2\nabla_{[\nu}\phi_{\rho]}=\nabla_{\nu}\phi_{\rho}-\nabla_{\rho}\phi_{\nu}$ and $2\nabla_{(\mu}\xi_{\nu)}=\nabla_{\mu}\xi_{\nu}+\nabla_{\nu}\xi_{\mu}$.. The parameter $\sigma\in\{+1,-1\}$ allows one to choose the sign in front of the free spin-1 Lagrangian. The value $\sigma=+1$ corresponds to the canonical sign of a unitary theory, while if $\sigma=-1$, the spin-1 sector will include negative-norm states. The free gauge transformations that leave $S_{0}$ invariant, i.e. $\delta_{0}S_{0}=0$, are
$\delta_{0}h_{\mu\nu}$ $=2\nabla_{(\mu}\xi_{\nu)}\,,\quad\delta_{0}A_{\mu}=\nabla_{\mu}\pi\;,$ (3)
$\delta_{0}\phi_{\mu}$ $=\nabla_{\mu}\epsilon+\frac{1}{2}i\lambda\gamma_{\mu}\epsilon\,,\quad\delta_{0}\bar{\phi}_{\mu}=\nabla_{\mu}\bar{\epsilon}+\frac{1}{2}i\lambda\bar{\epsilon}\gamma_{\mu}\,$, (4)
where $\xi_{\nu}$ is a vector gauge parameter, $\epsilon$ is a spinor gauge parameter and $\pi$ is a scalar one.
The equations of motion derived from the free Dirac gravitino Lagrangian in (2) describe the degrees of freedom of a gravitino gauge potential on a fixed de Sitter background. The physical mode solutions of these equations on global dS4 have been discussed in detail in [29, 34, 35], and it has been shown that they furnish a direct sum of fermionic discrete series unitary irreducible representations (UIRs) of $so(4,1)$. A curious field-theoretic feature of these representations is that physical modes of helicity 3/2-3/2 have a different positive-definite, $so(4,1)$-invariant inner product compared to the helicity +3/2+3/2 physical modes. In fact, the two inner products differ by a factor of 1-1. At the level of mode solutions, a different inner product can be used for each fixed-helicity sector (each fixed-helicity sector furnishes independently a discrete series UIR). However, from a quantum field theory viewpoint, the $so(4,1)$-invariant inner product must be the same for both helicities, if we insist on the locality of the theory [35]. Thus, the free-theory Dirac spin-3/2 Lagrangian used in (2) gives rise to indefinite norms upon quantisation. At the free level, there are two ways to achieve the unitary quantisation of the gravitino field on dS4. The first approach is to impose the on-shell anti-self-duality constraint to the gravitino field strength, which removes all negative-norm states, and thus quantise the chiral gravitino theory [35]. The second approach is to use an operator algebra perspective as a guiding principle, which leads to positive-definite inner product for the mode solutions that cannot be derived from a local Lagrangian [38].
We have shown that the gravitino Lagrangian in (2) is equivalent to the one used by Pilch, van Nieuwenhuizen, and Sohnius in [10], by properly expanding the Dirac spin-3/2 gravitino into a pair of symplectic Majorana gravitini, and performing a similarity transformation of the gamma matrices. We conclude that the gravitino sector of [10, 47] has indefinite norm states, i.e. the graviphoton kinetic term is not the only source of non-unitarity, which was not noticed in these two papers.
3 Consistent interactions at cubic order
Starting from the free theory (2), we classified all the possible local, cubic interactions among the various fields of the spectrum. We used the cohomological reformulation of the Noether method given in [32, 48], based on the antifield formalism of [49, 50, 51, 52, 53] and found two non-Abelian cubic vertices in the fermionic sector. First, there is the gravitational coupling of the Dirac spin-3/2 field, denoted $L_{1}^{(\phi-h)}$, with coupling constant $\alpha_{1}$. Second, we recover the electromagnetic coupling of the Dirac spin-3/2 field, denoted $L_{1}^{(\phi-A)}$, with coupling constant $\alpha_{2}$. These vertices, as well as the corresponding deformations to the gauge transformations, are given in Appendices A and B.
In the bosonic sector, there are two cubic vertices, the non-Abelian Einstein-Hilbert self-coupling $L_{1}^{EH}$, with coupling $\kappa$, and the Abelian gravitational coupling of the Maxwell field $L_{1}^{(h-A)}$, with coupling constant $\alpha_{3}$. Both of these couplings, as well as the corresponding deformations of the gauge transformations and of the gauge algebra, are given in e.g. [54] for both de Sitter and anti-de Sitter backgrounds.
We have shown that the Jacobi identities for the gauge algebra impose the following necessary conditions on the coupling constants:
$2\,\sigma\,\alpha_{2}^{2}+\alpha_{1}^{2}=0\;,\quad\alpha_{1}=-\frac{\kappa}{2}\;.$ (5)
As $\alpha_{2}$ and $\alpha_{1}$ are real constants, ensuring the reality of the deformed action $S$, the parameter $\sigma$ has to take the value 1-1. This leads to an opposite relative sign between the spin-1 and spin-2 kinetic terms, see eq. (2). Thus, we confirm the conclusions of [10] concerning the ghostly nature of the graviphoton, following a different path. Note that the derivation of eq. (5) does not require the explicit knowledge of the form of quartic vertices and of the corresponding gauge transformations.
4 The complete theory
4.1 Superalgebra, fields and gauge parameters
We denote $M_{ab}$ and $P_{a}$ $(a,b=0,1,2,3)$ the generators of $SO(4,1)$, the isometry group of 4-dimensional de Sitter spacetime. We introduce a supersymmetry charge, $Q$, which is a Dirac spinor, its conjugate $\bar{Q}$ and a $U(1)$ generator $T$. The non-vanishing commutators and anti-commutators are
$[M_{ab},M_{cd}]$ $=2\eta_{a[c}M_{d]b}-2\eta_{b[c}M_{d]a}\;,$ (6)
$\quad[P_{a},M_{bc}]$ $=2\eta_{a[b}P_{c]}\;,\quad[P_{a},P_{b}]=-\lambda^{2}M_{ab}\;,$
$[T,Q]$ $=i\frac{\lambda\kappa}{\sqrt{2}}Q\;,\quad[T,\bar{Q}]=-i\frac{\lambda\kappa}{\sqrt{2}}\bar{Q}\;,$ (7)
$[M_{ab},Q]$ $=-\frac{1}{2}\gamma_{ab}Q\;,\quad[M_{ab},\bar{Q}]=\frac{1}{2}\bar{Q}\gamma_{ab}\;,$ (8)
$[P_{a},Q]$ $=\frac{i\lambda}{2}\gamma_{a}Q\;,\quad[P_{a},\bar{Q}]=\frac{i\lambda}{2}\bar{Q}\gamma_{a}\;,$ (9)
$\{Q,\bar{Q}\}$ $=-\frac{1}{4}\gamma^{5}\gamma^{a}P_{a}+\frac{i\lambda}{8}\gamma^{5}\gamma^{ab}M_{ab}-\frac{1}{2\sqrt{2}\kappa}\gamma^{5}T\;,$ (10)
where the bosonic generators are taken to be anti-hermitian. We checked that the (super-)Jacobi identities are satisfied, and thus, the generators in (6)-(10) define a $\mathcal{N}=2$ de Sitter Lie superalgebra, with bosonic subalgebra isomorphic to $so(4,1)\oplus u(1)$. The de Sitter superalgebras have been classified in [55, 56, 45, 10].
One introduces a Grassmann-even gauge field $W_{\mu}$ and a Grassmann-even gauge parameter $\varepsilon$ that are both valued in the adjoint representation of the superalgebra:
$W_{\mu}:=e_{\mu}^{a}P_{a}+\tfrac{1}{2}\omega_{\mu}{}^{ab}M_{ab}+\bar{\psi}_{\mu}\,Q+\bar{Q}\,\psi_{\mu}+A_{\mu}T\;,$ (11)
$\varepsilon:=\varepsilon^{a}_{P}P_{a}+\frac{1}{2}\varepsilon_{M}^{ab}M_{ab}+\bar{\varepsilon}_{Q}\,Q+\bar{Q}\,\varepsilon_{Q}+\varepsilon_{T}\,T\;.$ (12)
The transformation of the gauge field $W_{\mu}$ under the gauge group is defined as $\delta^{G}\,W_{\mu}:=\partial_{\mu}\varepsilon+[W_{\mu},\varepsilon]\,.$ The gauge field component $e_{\mu}^{a}$ is assumed to be invertible, with inverse denoted $e^{\mu}_{a}$, and $e_{\mu}^{a}e^{\nu}_{a}=\delta^{\nu}_{\mu}$. As usual, $g_{\mu\nu}:=e_{\mu}^{a}\eta_{ab}e_{\nu}^{b}$. Following the gauge theory prescription, we introduce the curvatures of the gauge group
$\mathcal{R}_{\mu\nu}:=2\partial_{[\mu}W_{\nu]}+[W_{\mu},W_{\nu}]\;,$ (13)
$\mathcal{R}_{\mu\nu}=$ $\,\mathcal{R}_{\mu\nu}^{a}(P)\,P_{a}+\tfrac{1}{2}\mathcal{R}_{\mu\nu}{}^{ab}(M)M_{ab}+\bar{\mathcal{R}}_{\mu\nu}(Q)\,Q+\bar{Q}\,\mathcal{R}_{\mu\nu}(Q)+\mathcal{R}_{\mu\nu}(T)T\;.$ (14)
[…]
$\mathcal{R}_{\mu\nu}{}^{ab}(M)=R_{\mu\nu}{}^{ab}-2\lambda^{2}e_{[\mu}^{a}e_{\nu]}^{b}+\frac{i\lambda}{2}\bar{\psi}_{[\mu}\gamma^{5}\gamma^{ab}\psi_{\nu]}\;,$ (15)
where $R_{\mu\nu}{}^{ab}=2\partial_{[\mu}\omega_{\nu]}{}^{ab}+2\omega_{[\mu}{}^{cb}\omega_{\nu]}{}_{c}{}^{a}$ is the Riemann tensor. Also, $\mathcal{R}_{\mu\nu}(T)=F_{\mu\nu}-\frac{1}{\sqrt{2}\kappa}\bar{\psi}_{[\mu}\gamma^{5}\psi_{\nu]}\;$.
4.2 The action
Inspired by the construction of $\mathcal{N}=2$ supergravity in anti-de Sitter spacetime [57, 58] as presented in [59, 60], we propose the following action of $\mathcal{N}=2$ de Sitter supergravity:
$I[e,\omega,\bar{\psi},\psi,A]=$ $\int d^{4}x\Big(\;\frac{1}{32\kappa^{2}\lambda^{2}}\mathcal{R}_{\mu\nu}{}^{ab}(M)\mathcal{R}_{\rho\sigma}{}^{cd}(M)\varepsilon^{\mu\nu\rho\sigma}\varepsilon_{abcd}$ (16)
$-\frac{1}{8\kappa^{2}\lambda}\bar{\mathcal{R}}_{\mu\nu}(Q)\mathcal{R}_{\rho\sigma}(Q)\varepsilon^{\mu\nu\rho\sigma}+\frac{1}{4}\,e\,g^{\mu\rho}g^{\nu\sigma}\mathcal{R}_{\mu\nu}(T)\mathcal{R}_{\rho\sigma}(T)\Big)\;.$
We have shown that, [^3] Details of the calculation will be given in an extended version of this work. on the surface $\Sigma\equiv\frac{\delta I}{\delta\omega_{\mu}{}^{ab}}=0$, the action $I[e,\bar{\psi},\psi,A]=I|_{\Sigma}$ is invariant under the following local supersymmetry transformations,
$\delta e_{\mu}^{a}$ $=-\frac{1}{4}(\bar{\psi}_{\mu}\gamma^{5}\gamma^{a}\varepsilon_{Q}-\bar{\varepsilon}_{Q}\gamma^{5}\gamma^{a}\psi_{\mu})\;,$ (17)
$\delta\psi_{\mu}$ $=\mathcal{D}_{\mu}\varepsilon_{Q}+\frac{i\lambda}{2}\gamma_{\mu}\varepsilon_{Q}+\frac{\kappa}{4\sqrt{2}}\mathcal{R}_{\nu\rho}(T)\gamma^{\nu\rho}\gamma_{\mu}\varepsilon_{Q}\;,$ (18)
$\delta\bar{\psi}_{\mu}$ $=\mathcal{D}_{\mu}\bar{\varepsilon}_{Q}+\frac{i\lambda}{2}\bar{\varepsilon}_{Q}\gamma_{\mu}+\frac{\kappa}{4\sqrt{2}}\mathcal{R}_{\nu\rho}(T)\bar{\varepsilon}_{Q}\gamma_{\mu}\gamma^{\nu\rho}\;,$ (19)
$\delta A_{\mu}$ $=\frac{1}{2\sqrt{2}\kappa}(\bar{\varepsilon}_{Q}\gamma^{5}\psi_{\mu}-\bar{\psi}_{\mu}\gamma^{5}\varepsilon_{Q})\;,$ (20)
to all orders in the fermions, where
$\mathcal{D_{\mu}}\varepsilon_{Q}:=\left(\partial_{\mu}+\frac{1}{4}\omega_{\mu}{}^{ab}(e,\bar{\psi},\psi)\,\gamma_{ab}-i\frac{\lambda\kappa}{\sqrt{2}}A_{\mu}\right)\varepsilon_{Q}\;.$
The action above is the simplest de Sitter supergravity theory that can be constructed, and, to the best of our knowledge, it has never been given to all orders in the fermions before. It defines $\mathcal{N}=2$ de Sitter pure supergravity theory. To show its uniqueness, we rewrite it as follows, dropping the terms quartic in the fermions:
$I[e,\bar{\psi},\psi,A]=\int d^{4}x\,e\,\Bigg($ $\frac{1}{2\kappa^{2}}(R-2\Lambda)-\frac{1}{2\kappa^{2}}\bar{\psi}_{\mu}\gamma^{5}\gamma^{\mu\nu\rho}\mathcal{D}_{\nu}\psi_{\rho}$ (21)
$+\frac{i\lambda}{2\kappa^{2}}\bar{\psi}_{\mu}\gamma^{5}\gamma^{\mu\nu}\psi_{\nu}+\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$
$-\frac{1}{2\sqrt{2}\kappa}\bar{\psi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\psi_{\nu}\Bigg)+\mathcal{O}(\psi^{4})\,$,
where $\Lambda=3\lambda^{2}$ is the cosmological constant, $\mathcal{D}_{\mu}\psi_{\nu}=(\partial_{\mu}+\frac{1}{4}\omega_{\mu}{}^{ab}(e,\bar{\psi},\psi)\,\gamma_{ab}-i\frac{\lambda\kappa}{\sqrt{2}}A_{\mu})\psi_{\nu}$ , and $(*F)^{\mu\nu}=\frac{1}{2}e^{-1}\varepsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$. Now, it is easy to see that all the cubic terms appearing above exactly produce the exhaustive list of cubic couplings we have classified in the previous section. Following standard arguments, explained in e.g. the conclusion of Section 3 of [61], we conclude the uniqueness of the theory.
We have checked that, up to terms quartic in fermions and upon performing a similarity transformation of the gamma matrices, the action in eq. (21) coincides with the action written down in [10]. The important difference between our result and [10] is that we have explicitly constructed the terms quartic in fermions, and have shown gauge invariance of the action to all orders. As expected, the spin-1 field in (21) has the wrong sign in front of its kinetic term, making the appearance of negative-norm states unavoidable. However, as mentioned earlier, the graviphoton is not the only source of non-unitarity of the theory; the gravitino sector of the theory also contains negative-norm states.
5 Conclusion and Outlook
In this paper, we provided the complete action for $\mathcal{N}=2$ pure supergravity with positive cosmological constant and proved its uniqueness, as well as clarified that both the graviphoton and the gravitino give rise to negative-norm states from a quantum field theory perspective. We have thus extended the results of [10] in several directions. The complete action we found in this paper opens new directions in view of the renewed interest and new results concerning the Euclidean approach to de Sitter quantum gravity [36, 37, 38]. In particular, it is tempting to study the consistency of de Sitter supergravity in Euclidean signature, taking as our guide the sphere path integral [36, 37, 38]. Indeed, certain reality properties that are considered problematic in the Lorentzian setting, such as a complex action, turn out to be harmless in the Euclidean theory [38]. Quantum field theory in Lorentzian signature requires a real action for quantum mechanical consistency of the theory, but the real action for the gravitino (2) gives rise to an indefinite inner product [35]. However, this is not necessarily problematic as, ultimately, the gravitino should be viewed as part of gravitational system (gravitino coupled to gravity around dS4), and the unitarity of a quantum gravitational theory may be approached from a different perspective. Adopting a different perspective for unitarity should not come as a surprise considering the fact that de Sitter symmetries must annihilate all states in the Hilbert space already at the level of linearised gravity [62]. Having mentioned different perspectives, it was recently shown [38] that in the Euclidean path integral approach it is consistent for the gravitino action to be complex [38], see also [63]. In particular, a complex action for the gravitino on the four-sphere (Euclidean de Sitter) leads to 1-loop expressions for the sphere path integral that are not only real-valued but also are expressed in terms of unitary $so(4,1)$ characters [38], avoiding the problems of the Lorentzian theory. Therefore, it appears very promising to take the complete action of $\mathcal{N}=2$ de Sitter pure supergravity that we have established in this paper, find its counterpart in Euclidean signature and study its partition function, as was done in the 2D case [46]. This is something we leave for future work, as well as the non-linear completion of the free unitary chiral supergravity multiplet found in [35]. Finally, our results concerning the explicit knowledge of the complete set of interaction vertices of the simplest, pure de Sitter supergravity, may also be useful for the determination of higher-point cosmological correlators for fermions, see [64] for a related recent work.
Addressing these questions will contribute to the ongoing effort to obtain a theoretical dataset for quantities central to (Euclidean) quantum gravity with positive cosmological constant.
Acknowledgments
We would like to thank Atsushi Higuchi for many useful discussions and for sharing his helpful notes. We would also like to thank Dionysios Anninos, Evgeny Skvortsov and Philippe Spindel for useful discussions. The packages xTras [65] and FieldsX [66] of the xAct suite of Mathematica packages for tensor computer algebra were used in several computations. S.T. is a research fellow of the F.R.S.-FNRS supported by the ASP fellowship FC 54793 MassHighSpin. The work of V. A. L. is supported by the ULYSSE Incentive Grant for Mobility in Scientific Research [MISU] F.6003.24, F.R.S.-FNRS, Belgium. V. A. L. thanks the Department of Mathematics of the University of York (UK) for hospitality where part of his work was carried out. The work of N. B. was partly supported by the F.R.S.-FNRS PDR grant number T.0047.24.
Appendix A Gravitational coupling of the Dirac spin-3/2 field
The non-Abelian cubic interaction term involving a the massless spin-2 field $h_{\mu\nu}$ and a massless Dirac spin-3/2 field, $\phi_{\mu}$, is $S_{1}^{(\phi-h)}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha_{1}L_{1}^{(\phi-h)}$, with
$L_{1}^{(\phi-h)}=$ $\;\bar{\phi}_{\alpha}\gamma^{\alpha\beta\gamma}\gamma^{\mu\nu}\gamma^{5}\phi_{\gamma}\nabla_{\mu}h_{\nu\beta}+2\bar{\phi}_{\delta}\gamma^{\gamma\delta\alpha}\gamma^{5}\nabla_{\gamma}\phi_{\alpha}h$ (22)
$+2\bar{\phi}_{\alpha}\gamma^{\gamma\delta\beta}\gamma^{5}\nabla_{\delta}\phi_{\beta}h^{\alpha}{}_{\gamma}+2\bar{\phi}_{\delta}\gamma^{\gamma\delta\alpha}\gamma^{5}\nabla_{\alpha}\phi_{\beta}h^{\beta}{}_{\gamma}$
$-2\bar{\phi}_{\delta}\gamma^{\gamma\delta\alpha}\gamma^{5}\nabla_{\beta}\phi_{\alpha}h^{\beta}{}_{\gamma}-2i\lambda\bar{\phi}_{\beta}\gamma^{\alpha\beta}\gamma^{5}\phi_{\gamma}h_{\alpha}{}^{\gamma}$
$+2i\lambda\bar{\phi}_{\gamma}\gamma^{\alpha\beta}\gamma^{5}\phi_{\beta}h_{\alpha}{}^{\gamma}-2i\lambda\bar{\phi}_{\alpha}\gamma^{\alpha\beta}\gamma^{5}\phi_{\beta}h\;.$
The corresponding deformations of the gauge transformations are
$\delta^{(\phi-h)}_{1}h_{\mu\nu}$ $=\alpha_{1}\Bigl(\bar{\epsilon}\gamma_{(\mu}\gamma_{5}\phi_{\nu)}-\bar{\phi}_{(\mu}\gamma_{\nu)}\gamma_{5}\epsilon\Bigr)\,$, (23)
$\delta^{(\phi-h)}_{1}\phi_{\mu}$ $=\alpha_{1}\Bigl(\gamma^{\alpha\beta}\epsilon\,\nabla_{\alpha}h_{\beta\mu}-i\lambda\gamma^{\nu}\epsilon\,h_{\mu\nu}+2i\lambda\gamma_{\nu}\phi_{\mu}\xi^{\nu}-\gamma_{\alpha\beta}\phi_{\mu}\nabla^{\alpha}\xi^{\beta}+8\Phi_{\mu\nu}\xi^{\nu}\Bigr)\,$,
$\delta^{(\phi-h)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(\phi-h)}_{1}\phi_{\mu}$ , and $\delta^{(\phi-h)}_{1}A_{\mu}=0$ .
Appendix B Electromagnetic coupling of the Dirac spin-3/2 field
The non-Abelian coupling between the massless spin-1 field and the Dirac spin-3/2 is $S^{(\phi-A)}_{1}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha_{2}L^{(\phi-A)}_{1}$ , with
$L^{(\phi-A)}_{1}$=  $2\Bigl(\bar{\phi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\phi_{\nu}+i\lambda A_{\alpha}\bar{\phi}_{\gamma}\gamma^{\alpha\beta\gamma}\gamma^{5}\phi_{\beta}\Bigr)\;.$ (24)
The corresponding deformations to the gauge transformations are
$\delta^{(\phi-A)}_{1}\phi_{\mu}$ $=\alpha_{2}\Bigl(-2i\lambda\phi_{\mu}\pi+2i\lambda A_{\mu}\epsilon-\frac{1}{2}F_{\alpha\beta}\gamma^{\alpha\beta}\gamma_{\mu}\epsilon\Bigr)\,$,
$\delta^{(\phi-A)}_{1}A_{\mu}$ $=2\,\alpha_{2}\,\sigma\Bigl(\bar{\epsilon}\gamma^{5}\phi_{\mu}-\bar{\phi}_{\mu}\gamma^{5}\epsilon\Bigr)\,$, (25)
$\delta^{(\phi-A)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(\phi-A)}_{1}\phi_{\mu}$ , and $\delta^{(\phi-A)}_{1}h_{\mu\nu}=0$ .